camille sikorski was divorced last year. she currently provides a home for her 15 year old daughter kaly. kaly lived in camille's home for the entire year, and camille paid for all the costs of maintaining the home. she received a salary of $70,000 and contributed $4,800 of it to a qualified retirement account (a for AGI deduction). she also received $9,000 of alimony from her former husband. finally, camille paid $15,900 of expenditures that qualified as itemized deductions.
a) what is camille's taxable income?

(1) gross income
(2) for AGI deductions
(3) adjusted gross income
(4) standard deduction
(5) itemized deductions
(6)
   taxable income
can you please type it out thank you
Answer #1

1. Gross income is the amount of money we earn before anything is taken out for taxes or other deductions.

GI = 70000 (Salary) + 9000 (Alimony) = $ 79000

2. Adjusted gross income (AGI) is a measure of income calculated from your gross income and used to determine how much of your income is taxable.

AGI Deduction = 4800

3. Adjusted gross income (AGI) is an individual's total gross income minus specific deductions.

AGI = Gross Income - AGI Deduction - Itemized Deduction

   = 79000 - 4800 - 15900

= $ 58300

4. Standard deduction is essentially a flat amount subtracted from the salary income before calculation of taxable income.

Standard Deduction = $ 4800

5.  An itemized deduction is an expenditure on eligible products, services, or contributions that can be subtracted from adjusted gross income (AGI) to reduce your tax bill.

Itemized deduction = $ 15900

6.  Camille Taxable Income is computed as follows

Taxable Income = Gross Income - standard deduction - itemized deduction

= 70000 - 4800 - 15900

= $ 49300

NOTE:  Alimony is payable only when the couple is legally divorced. Alimony is not taxable if: It is paid in lump-sum or a one-time receipt in the form of cash.